# Shadow in the Cache: Unveiling and Mitigating Privacy Risks of KV-cache in LLM Inference

### v1

@@ III-B2 KV-cache Collision Attack
Unlike the KV-cache Inversion Attack, the KV-cache Collision Attack operates by iteratively generating KV-cache entries for candidate tokens from the model’s vocabulary. For each candidate token appended to a partially reconstructed input sequence, the attack performs local model inference, generates the corresponding KV-cache, and compares it against the intercepted KV-cache to determine if the candidate token is correct for the current position.
Fundamental Principle of the Collision Attack. The collision attack reconstructs user input by identifying the token at each position that minimizes a distance metric (e.g., L2 distance) between a locally generated KV-cache and the leaked KV-cache. Its effectiveness hinges on the assumption that the distance for the correct token, $dis_{target}$, is statistically separable from the distances of all other incorrect tokens from the vocabulary, $dis_{other}$, allowing for its unambiguous identification.
General Attack Procedure. As illustrated in Figure 3b, the attack reconstructs the user’s input token-by-token. After initially intercepting the target KV-cache from the cloud service, the attacker enters an iterative loop for each unknown token position. First, guided by the already-reconstructed prefix, the attacker’s local model instance predicts a probabilistically-ranked list of candidate tokens. Next, for each high-probability candidate, the attacker performs a forward pass to generate a local KV-cache segment and compares it against the corresponding segment of the target KV-cache to test for a match (a “collision”). Upon a successful match, the correct token is confirmed and appended to the reconstructed prefix, and the attacker proceeds to the next position. This process continues until the entire input sequence is recovered.
Implementation Optimizations for the Collision Attack. A naive collision attack that processes the entire model vocabulary is impractical due to prohibitive VRAM consumption and execution time. We therefore introduce some synergistic optimizations to make the attack practical. First, instead of analyzing all prioritized candidates, we apply batched statistical validation, testing a subset of candidate tokens and computing the distribution of their distances ($dis_{batch}$) from the leaked KV-cache. The correct token is then identified as any distinct statistical outlier within this batch, under the assumption that the batch provides a representative sample of the distance distribution for incorrect tokens. Second, we employ a probability-guided search. Using the model’s own predictions based on the already reconstructed prefix, we prioritize collision tests for high-likelihood candidate tokens. Furthermore, we prune the search space by truncating the vocabulary and discarding tokens with very low probabilities. Together, these techniques reduce the memory footprint and bring the attack latency down to the order of seconds.
Enhancing Collision Attack with Prior Knowledge. A key challenge in the collision attack is setting an optimal classification threshold to distinguish the correct token’s distance ($dis_{target}$) from the distribution of incorrect token distances ($dis_{other}$). A basic attacker, lacking ground-truth data, must rely on a static, heuristic threshold, typically derived from the statistics of observed incorrect token distances (e.g., $\mu_{other}-3\sigma_{other}$). However, an adversary capable of performing chosen-plaintext attacks can generate KV-caches for known inputs. This prior knowledge enables a more precise statistical characterization of both the $dis_{target}$ and $dis_{other}$ distributions, allowing for a more effective, adaptive threshold that significantly boosts reconstruction accuracy.
The need for an adaptive threshold is formalized by the success probability for a token at rank $r$ in the probability-sorted vocabulary, as defined in Eq. (5):
$P(\text{success}|t)=P(dis_{other}>t)^{r-1}\times P(dis_{target}<t).$ (5)
This equation demonstrates that the optimal threshold $t$ is dependent on the token’s expected rank $r$, as it must balance the trade-off between correctly rejecting the r1r-1 preceding (incorrect) candidates and accepting the true token. A static heuristic, by contrast, cannot effectively manage this rank-dependent trade-off.
The KV-cache Collision Attack overcomes the limitations of the Inversion Attack concerning specific KV-cache and model attention mechanism requirements. An attacker who intercepts KV-cache from potentially any layer of various LLM architectures can, in principle, employ this collision-based approach to reconstruct user inputs. The practical effectiveness and performance of this attack will be demonstrated and evaluated in subsequent experimental sections.
@@ III-B3 KV-cache Injection Attack
The KV-cache Injection Attack leverages the LLM’s own semantic capabilities to exfiltrate information, operating differently from direct state-matching methods. The KV-cache contains substantial information from user-model interactions, however, itself (comprising only $k$ and $v$ vectors for past tokens) lacks the $q$ vectors for the current generation step and thus cannot be directly used to continue model inference without a new input token. The core principle involves an attacker appending a crafted instruction (e.g., “Repeat the previous content.”) to an intercepted KV-cache, as showed in Figure 3c. This action provides the necessary query, key, and value vectors to resume inference from the compromised state, effectively tricking a standard LLM into executing the command within the context of the user’s private data. This can compel the model to “echo” or regenerate the sensitive input stored implicitly within the cache. While this method is generally less precise than the collision attack, its primary advantage is speed, requiring only a single forward inference pass. The viability of this attack vector reveals a crucial requirement for any comprehensive defense: a protected KV-cache must be rendered semantically unintelligible to prevent an unauthorized LLM from meaningfully parsing its contents or continuing generation from it.
We conduct comprehensive evaluations on these three attacks in Section V. The experimental results demonstrate the practicality and effectiveness of input reconstruction attacks from KV-cache, underscoring the urgent need for developing an effective defense technique.
IV KV-Cloak: A Defense Mechanism for
KV-cache Obfuscation
IV-A Motivation for KV-Cloak
Limitations of Existing Privacy-Preserving Techniques. While several techniques exist for data privacy, they prove impractical or inadequate for securing the LLM KV-cache, highlighting the need for a specialized solution. The introduction to existing privacy-preserving techniques and their limitations is as follows.
  • Cryptographic Methods: Standard cryptographic techniques [4, 31], such as symmetric AES encryption, provide confidentiality by ensuring data remains in ciphertext during storage and transmission, requiring decryption only for legitimate use. A more advanced approach, Homomorphic Encryption (HE) [22, 2], allows for direct computation on encrypted data, theoretically enabling parts of the attention mechanism to operate on the KV-cache without ever decrypting it. However, despite their strong security guarantees, the immense computational overhead and latency introduced by these methods are prohibitive. Given that the KV-cache can be tens or hundreds of gigabytes, applying full encryption or HE is unsuitable for the high-throughput, low-latency requirements of LLM inference.
  • Differential Privacy (DP): Differential Privacy is a rigorous model that protects individual privacy by injecting carefully calibrated random noise into data or query results [1, 16, 39]. The fundamental goal is to make the output of a computation statistically insensitive to the presence or absence of any single individual’s data in the input dataset. In the context of the KV-cache, this would involve adding noise to the Key and Value vectors. The primary limitation of this approach is the inherent trade-off between privacy and utility. To achieve a meaningful level of privacy, the required amount of noise would significantly degrade the LLM’s inference accuracy to an unacceptable degree.
  • KV-Shield: KV-Shield [38] is a lightweight obfuscation scheme and, to our best knowledge, the only privacy-preserving method designed specifically for the KV-cache. Its process involves synchronously permuting the rows of the model’s attention layer parameters ($W_{q},W_{k},W_{v}$). This operation effectively shuffles the element positions within the resulting KV-cache vectors, obfuscating them before an attention score is calculated. The obfuscated attention output is then de-obfuscated for subsequent steps. However, this approach has critical flaws. From a security perspective, its fixed shuffling pattern does not alter the statistical properties of the data, leaving it vulnerable to collision attacks. For compatibility, its technique is incompatible with modern LLM architectures that utilize features like Rotary Positional Embedding (RoPE), as the shuffling interferes with the position-dependent calculations.
Design Objectives. The shortcomings of existing methods reveal an urgent need for a defense mechanism that is both secure and practical. To be effective, such a solution must achieve three key objectives:
  • High Security: It must provide robust protection against targeted reconstruction attacks.
  • Minimal Accuracy Impact: It must operate without causing any degradation in the model’s output accuracy.
  • Low Performance Overhead: It must introduce only negligible latency into the inference pipeline.
To meet these challenges, we propose KV-Cloak, a novel defense mechanism designed to provide strong, lossless, and efficient protection for the KV-cache.
IV-B A Naive Defense: Obfuscation via Reversible Linear Transforms
We first evaluate a baseline defense that employs a reversible linear transform to obscure the statistical properties of the $k$ and $v$ vectors. In the context of modern inference frameworks like vLLM, we represent all key vectors for a single attention head within a KV-cache block as a matrix $K\in\mathbb{R}^{b\times d}$, where $b$ is the number of vectors per block and $d$ is the attention head dimension. As the obfuscation method is analogous for both keys and values, we use $K$ as our canonical example. The transformation is defined as:
$K^{\prime}=SKM,$ (6)
where $S\in\mathbb{R}^{b\times b}$ and $M\in\mathbb{R}^{d\times d}$ are secret, randomly generated, and invertible matrices. The objective is to transform the original matrix $K$ without altering its dimensions, preventing an adversary from directly recovering its contents.
Vulnerability Analysis. Despite its simplicity, this linear transform is fundamentally insecure under a Chosen-Plaintext Attack (CPA). While an adversary cannot craft an arbitrary matrix $K$ from scratch—as its contents are determined by the model’s embedding and projection layers—they can still exert significant influence through carefully constructed inputs. This capability is sufficient to break the scheme.
The fixed linear nature of the transformation makes it highly susceptible to algebraic cryptanalysis. An adversary can mount a differential attack by choosing two plaintexts, $K_{1}$ and $K_{2}$, thereby controlling the difference $\Delta K=K_{1}-K_{2}$ and observing the corresponding output difference $\Delta K^{\prime}=S(\Delta K)M$. By systematically choosing $\Delta K$ to be a series of standard basis matrices (i.e., matrices with a single non-zero entry), the adversary can isolate the columns of $S$ and the rows of $M$. This process allows for the full recovery of the secret matrices $S$ and $M$ (up to a trivial scaling ambiguity), leading to a complete compromise of the scheme.
IV-C The Improved Defense: One-Time Pad Block-wise Shuffling
Eliminating Redundant Positional Information in KV-cache. To enhance the security of the obfuscation scheme, we introduce additional randomness by eliminating redundant positional information inherent to KV-cache storage. During attention mechanism computations, the physical ordering of $k,v$ vectors in the KV-cache is an unnecessary redundancy for inference. In reality, positional information critical for computation is already embedded into the elements of $q$ and $k$ vectors through rotary positional encoding.
Figure 4: Shuffling $k,v$ vectors within a KV-cache block preserves attention computation results while disrupting positional order.
By randomly permuting $k,v$ within each block while maintaining their correspondence to input tokens $x$, we eliminate positional clues that attackers could exploit, and the results of the attention mechanism were not affected (as Figure 4). This permutation increases the computational complexity for attackers to match plaintext-ciphertext pairs by a factor of $b!$, where $b$ denotes the block size (default $b=16$ in vLLM).
The enhanced obfuscation process employs a one-time pad permutation matrix $\hat{P}$ to reorder vectors within blocks before applying invertible matrices $S$ and $M$:
$K^{\prime}=S\hat{P}KM.$ (7)
Crucially, $\hat{P}$ does not need to be stored as a key since the deobfuscated $\hat{P}K$ can directly participate in subsequent computations without reconstruction. This design choice preserves computational efficiency.
Additive Mask Matrices with Position-Identifiable Features. However, when the original $K$ matrix has rank $1$ (i.e., all row vectors are identical), the entropy of the permutation $\hat{P}$ will collapse. To prevent rank deficiency, KV-Cloak introduce an additive masking matrix $A$ before permutation. Thereby, we can ensure the transformed matrix $(K+A)$ maintains sufficient rank to preserve cryptographic strength of $\hat{P}$.
Since the one-time pad permutation matrix $\hat{P}$ is not stored, computing $\hat{P}A$ during inverse obfuscation becomes infeasible, and thus $\hat{P}K=\hat{P}(K+A)-\hat{P}A$ cannot be derived from $\hat{P}(K+A)$. To avoid storing the one-time pad permutation matrix $\hat{P}$, we exploit the sparsity and small-elements properties of $k$ and $v$ vectors to design the additive matrix $A$ with additional structural constraints. This enables us to infer $\hat{P}A$ based on the data characteristics of $\hat{P}(K+A)$.
Figure 5: Using an excerpt from “The Bitter Lesson” [30] (see Appendix -B) as the input of LLaMA-3.2-1b, the KV-cache data element distribution generated during the inference process.
As shown in Figure 5, we observe that the elements of $k$ and $v$ vectors typically remain below a small threshold $\theta_{K}$ (e.g., 100). This allows us to construct a matrix $A$ where positional information is embedded into $K+A$ via additive operations using values significantly larger than $\theta_{K}$ in specific row vectors. After applying row permutations, the positional identifiers in $\hat{P}(K+A)$ remain identifiable. During inverse obfuscation, the positional information of $\hat{P}A$ can be directly identified from $\hat{P}(K+A)$ by detecting these outliers. Subtracting $\hat{P}A$ from $\hat{P}(K+A)$ yields $\hat{P}K$ for subsequent inference, drastically reducing storage overhead for the one-time pad key.
The complete obfuscation transformation, as shown in Eq. (8), incorporates additional matrices $\hat{P}$ and $A$ alongside invertible linear transformations $S$ and $M$:
$K^{\prime}=S\hat{P}(K+A)M,$ (8)
[…]
IV-D Implicit Obfuscation via Operator Fusion
To mitigate the performance overhead introduced by the obfuscation and de-obfuscation processes during runtime, we now demonstrate how the obfuscation operators can be fused directly into the model’s weight matrices. This allows the primary cost of obfuscation to be paid offline, resulting in a minimal impact on online inference latency.
Fusing the Obfuscation Operator into Model Weights. We define three transformed vectors $(q^{m},k^{m},v^{m})$ and a transformed output projection matrix $W_{o}^{m}$ using two secret, invertible matrices, $M_{1}$ and $M_{2}$. Let $q,k,v$ be the original query, key, and value vectors, and let $W_{o}$ be the original output projection matrix. The transformations are defined as follows:
These transformations are specifically designed to leave the core attention mechanism computations invariant. The attention scores remain unchanged, as shown by the dot product:
$=q_{i}(M_{1}M_{1}^{-1})^{\top}k_{j}^{\top}=q_{i}k_{j}^{\top}.$
Similarly, the output of the attention head after projection remains identical:
$=v_{j}M_{2}(M_{2}^{-1})W_{o}^{\top}=v_{j}W_{o}^{\top}.$
Next, we incorporate the projection from the input embedding $x$ and the application of RoPE, denoted by the position-dependent matrix $R_{\Theta,i}^{d}$. The equations for the transformed vectors become:
From these formulations, we can observe the path to operator fusion. The secret matrices $M_{2}$ and $M_{2}^{-1}$, which obfuscate the value vectors, can be straightforwardly fused into the model’s weights by pre-computing new weight matrices $(W_{v}^{m})^{\top}=W_{v}^{\top}M_{2}$ and $W_{o}^{m}=W_{o}(M_{2}^{-1})^{\top}$.
However, fusing $M_{1}$ and $M_{1}^{-1}$ into $W_{k}$ and $W_{q}$ is more complex due to the intermediate application of the position-dependent RoPE matrix $R_{\Theta,i}^{d}$. The fusion becomes feasible under a specific algebraic condition: if the secret matrix $M_{1}$ and the RoPE matrix $R_{\Theta,i}^{d}$ commute, i.e., $R_{\Theta,i}^{d}M_{1}=M_{1}R_{\Theta,i}^{d}$. If this condition holds, we can rewrite the expression for $k^{m}$:
$k^{m}=xW_{k}^{\top}R_{\Theta,i}^{d}M_{1}=x(W_{k}^{\top}M_{1})R_{\Theta,i}^{d}.$ (13)
This allows us to pre-compute a new, position-independent weight matrix $(W_{k}^{m})^{\top}=W_{k}^{\top}M_{1}$. A similar argument applies to fusing $(M_{1}^{-1})^{\top}$ into $W_{q}$.
Therefore, the viability of this operator fusion approach hinges on our ability to construct a secret matrix $M_{1}$ that commutes with the RoPE matrices. In Appendix -A, we formally prove that this commutativity is achieved when $M_{1}$ is structured as a block-diagonal matrix composed of $2\times 2$ rotation-scaling sub-matrices, a structure analogous to that of RoPE itself. Furthermore, to ensure the reversibility of the entire transformation for lossless de-obfuscation, $M_{1}$ must be constructed to be invertible.
By constraining $M_{2}$ to also be a random, invertible rotation-scaling matrix, and by carefully controlling the range of rotation-scaling coefficients in both $M_{1}$ and $M_{2}$, we ensure that the transformed padding tokens (which involve the affine transformation with matrix $A$) remain identifiable as outliers. These conditions allow the obfuscation operators ($M_{1},M_{2}$) and their corresponding de-obfuscation counterparts ($M_{1}^{-1},M_{2}^{-1}$) to be fused into the attention layer parameters $W_{q},W_{k},W_{v},\text{and }W_{o}$. Let the model parameters after operator fusion be denoted by $W_{q}^{m},W_{k}^{m},W_{v}^{m},W_{o}^{m}$. The new, fused weights are pre-computed as follows:
By deploying the LLM with these pre-computed fused weights, we effectively integrate the random invertible block-diagonal matrix $M_{1}$ and the general random invertible matrix $M_{2}\in\mathbb{R}^{d\times d}$ into the attention layer parameters. This eliminates the need for explicit online matrix multiplications with $M_{1}$ and $M_{2}$, significantly reducing performance overhead.
Ultimately, the KV-Cloak obfuscation formula, leveraging operator fusion, is applied as follows:
$K^{\prime}=S\hat{P}(K^{m}+A),$ (15)
where $K^{m}=KM$ represents the KV-cache generated by the inference process using the fused model weights. It is crucial to control the range of the rotation and scaling factors within the matrix $M$ to ensure that the additive mask $A$ can still reliably embed its positional information via the outlier mechanism.
@@ IV-D Implicit Obfuscation via Operator Fusion
TABLE I: Effectiveness of the inversion, collision, and injection attacks against the unprotected KV-cache from different model layers. Here, “collision+” denotes the collision attack enhanced with prior knowledge.
Model Metric Inversion Collision Collision+ Injection
@@
LLaMA-7B BERTScore ($\uparrow$) 1.000 0.065 0.092 0.449 0.769 0.611 1.000 1.000 1.000 0.765
@@
LLaMA-3.1-8B-Distilled BERTScore ($\uparrow$) 0.083 0.062 0.081 0.642 0.492 0.635 0.894 0.258 0.762 0.610
ROUGE-L ($\uparrow$) 0.000 0.000 0.001 0.633 0.227 0.413 0.868 0.122 0.479 0.421
@@
With this optimization, the online computational overhead is drastically reduced. The matrix multiplication with $M$ is now part of the offline weight preparation. The online cost for obfuscating and de-obfuscating one KV-cache block is now dominated by the multiplications with $S,\hat{P},S^{-1}$, totaling approximately $b^{3}+2b^{2}d$ floating-point operations. For the LLaMA-3.1-8B example ($b=16,d=128,D=4096$), this overhead is merely $0.83\%$ of the cost of re-computing the KV-cache block, rendering the performance impact minimal. Detailed analysis can be found in Appendix -F.
V Evaluation
V-A Experimental Setup
Models. We first select several state-of-the-art models from the popular LLaMA [9] and Qwen [37] series, including LLaMA-3.2-1B, LLaMA-3.2-3B-Instruct, LLaMA-3.1-8B, Qwen2.5-Math-7B, and DeepSeek-R1-Distill-LLaMA-8B (LLaMA-3.1-8B-Distilled) [10], which is fine-tuned from LLaMA-3.1-8B. In addition to these five models that use Grouped-Query Attention (GQA), we also leverage a classic models with Multi-Head Attention (MHA), namely LLaMA-7B [32], to demonstrate the generalizability of our method.
Datasets. To evaluate the effectiveness of our attacks, we use the lmsys-chat-1m dataset [43], which is collected from a real-world inference service and contains a rich variety of dialogues and instructions. We randomly sample 1,000 instances from this dataset as the test set for validating attack effectiveness. To assess the impact of KV-Cloak on model accuracy and inference latency, we employ standard language model benchmarks, including MMLU [13, 12] and SQuAD [25].
Attack and Defense Evaluation Metrics. BERTScore [41] and ROUGE-L [19] are used to measure the similarity between the original input and the text reconstructed from the KV-cache. BERTScore, which is based on the all-mpnet-base-v2 model, is better at capturing semantic similarity, while ROUGE-L primarily reflects lexical-level precision and recall.
V-B Attack Effectiveness
In this section, we first evaluate the privacy leakage risk of using the KV-cache by implementing our proposed inversion, collision, and injection attacks. Additional experiments, such as the ablation study of attacks, can be found in Appendix -C.
Experimental Settings. We attack 7 LLMs as described in Section V-A. Notably, while attacking the LLaMA-3.1-8B-Distilled, which is a fine-tuned version, we assume that the attacker can only get access to its base model (i.e., LLaMA-3.1-8B). It can be utilized to validate the attack effectiveness against fine-tuning models. Furthermore, the inversion and collision attacks require KV-cache data from only a single layer. Given that models have varying numbers of layers, we select the KV-cache from the first, middle, and last layers to perform the inversion and collision attacks. In contrast, the injection attack is conducted on the complete KV-cache data, encompassing all layers.
Results of KV-cache Inversion Attack. As shown in Table I, our experimental results generally align with our analysis: a successful inversion attack requires two sufficient conditions to be met: (1) access to the KV-cache from the model’s first layer, and (2) an invertible attention projection matrix, which is characteristic of the MHA mechanism. The inversion attack can reconstruct the user input with near-perfect fidelity (approaching 100%) from the first-layer KV-cache of MHA-based models. In contrast, reconstruction attempts using the KV-cache from subsequent layers yield completely unintelligible results (mostly <0.1<0.1).
However, we also find that for the LLaMA-3.2-1B model using GQA, the inversion attack achieved a reconstruction fidelity of nearly 100% on the first-layer KV-cache. This may be because the ratio of the rank of this model’s attention projection matrix to its hidden state dimension is particularly high. Consequently, even though the projection matrix is not a square, invertible matrix, the least squares method can still recover the input to the attention layer with high accuracy.

### v3

@@ III-B2 KV-cache Collision Attack
Unlike the Inversion Attack which suffers from architectural constraints, the KV-cache Collision Attack is a universal, forward-matching adversary applicable to any layer. Instead of algebraically reversing the projection, this attack reframes input reconstruction as a search problem: identifying the token in the vocabulary that, when processed by a local model, yields a KV-cache entry most similar to the intercepted target.
Fundamental Principle: Reconstruction as Optimization. The core premise is that the KV-cache exhibits distance-preserving properties for identical tokens under the same context. We formalize the attack as finding the optimal token $t^{*}$ at position $i$ that minimizes the distance metric $\mathcal{D}$ (e.g., Frobenius norm) between the leaked cache $K_{\text{leaked}}$ and the local generation $K_{\text{local}}(t)$:
$t^{*}_{i}=\mathop{\arg\min}_{t\in\mathcal{V}}\mathcal{D}\left(K_{\text{leaked}}^{(i)},K_{\text{local}}^{(i)}(t|x_{<i})\right).$ (5)
The attack’s success hinges on statistical separability: the distance for the ground-truth token, $d_{\text{target}}$, must be statistically distinguishable from the distribution of distances for incorrect tokens, $d_{\text{other}}$.
Attack Procedure. As illustrated in Figure 2b, the attack proceeds iteratively token-by-token. For each unknown position $i$: (1) Candidate Generation: The adversary selects candidate tokens from the model’s vocabulary $\mathcal{V}$ to test. (2) Local Simulation: For each candidate token $t$, the adversary performs a forward pass using their local model instance to generate the corresponding KV-cache entry $K_{\text{local}}^{(i)}(t)$. (3) Collision Matching: The adversary computes the distance $\mathcal{D}$ between the local entry and the intercepted target $K_{\text{leaked}}^{(i)}$. The candidate that minimizes this distance (i.e., creates a “collision” with the leaked state) is identified as the correct token $x_{i}$, appended to the sequence, and the process advances to $i+1$.
Implementation Optimizations. A naive exhaustive search over the entire vocabulary ($\mathcal{V}\approx 10^{5}$) entails prohibitive latency and VRAM consumption, as verifying the global minimum distance requires inferring every candidate token. To render the attack practical, we implement three synergistic optimizations that transform the problem from a global search into an efficient sequential decision process: (1) Batched Outlier Detection (Enabling Early Exit). Instead of calculating distances for the full vocabulary to find the minimum, we process candidates in batches. We fundamentally shift the decision logic: rather than comparing a candidate against all other tokens in $\mathcal{V}$, we determine if a candidate constitutes a statistical outlier within its current batch (i.e., $d_{t}<\mu_{\text{batch}}-3\sigma_{\text{batch}}$). Crucially, this allows for an early exit: once a collision is statistically confirmed within a batch, the search terminates immediately without evaluating the remaining vocabulary. (2) Probability-Guided Prioritization. To maximize the efficacy of the early exit mechanism, we employ a probability-guided search. We sort the candidate tokens based on the model’s predicted probabilities $P(x_{i}|x_{<i})$. This reordering ensures that the correct token—which typically carries a high probability—is placed in the first few batches. Consequently, the collision is triggered near the start of the search process, drastically reducing the average number of inferences required. (3) Search Space Pruning. Distinct from prioritization, we explicitly truncate the search space by discarding the “long tail” of tokens with negligible probabilities (e.g., keeping only the top-$k\%$ candidates). While prioritization accelerates the average case, pruning bounds the worst-case complexity. It prevents the adversary from wasting resources iterating through tens of thousands of implausible tokens in the rare event that the target is not found in the high-probability regions.
Enhancement via Chosen-Plaintext Attack (CPA). A critical challenge is determining the optimal decision threshold $\tau$. A static heuristic (e.g., $3\sigma$-rule) is often suboptimal due to the varying entropy of different tokens. However, an adversary with CPA capabilities—who can feed known inputs to the model and observe the resulting KV-cache—can profile the precise statistical behavior of $d_{\text{target}}$ and $d_{\text{other}}$.
This prior knowledge allows for an adaptive thresholding strategy. We formulate the success probability for a candidate at rank $r$ as:
$P(\text{success}|r)=P(d_{\text{other}}>\tau)^{r-1}\times P(d_{\text{target}}<\tau).$ (6)
This formulation explicitly captures the trade-off: a stricter threshold minimizes False Positives (accepting an incorrect predecessor) but risks False Negatives (rejecting the target). By dynamically adjusting $\tau$ based on the expected rank $r$ and the profiled distributions from CPA, the adversary maximizes $P(\text{success})$, significantly boosting reconstruction fidelity compared to static baselines.
@@ III-B3 KV-cache Injection Attack
Unlike Inversion and Collision attacks which rely on precise algebraic state matching, the KV-cache Injection Attack exploits the semantic instruction-following capabilities of LLMs to exfiltrate private contexts.
Architectural Constraint: The Missing Query. Architecturally, a stand-alone KV-cache is computationally dormant. The self-attention mechanism requires a current Query vector ($Q$) to drive the attention scoring ($\text{Softmax}(QK^{\top})$) and state transitions. Consequently, an adversary cannot simply “resume” inference from a stolen cache; a new input stimulus is required to generate the necessary activation state.
Attack Mechanism: Contextual Hijacking. To overcome this, we weaponize the LLM’s instruction-following nature. As shown in Figure 2c, the adversary appends a crafted directive (e.g., “Repeat the previous content.”) to the intercepted KV-cache. This injection serves two purposes: (1) Stimulus Generation. The directive creates the requisite $Q$ vectors to activate the attention mechanism. (2) Forced Exfiltration. The LLM attends to the stolen $K/V$ pairs as historical context. Bound by its alignment training, the model executes the instruction, effectively “echoing” or summarizing the latent private information stored in the cache.
Strategic Advantages: Robustness and Efficiency. This attack offers distinct advantages in robustness compared to the Collision Attack. Specifically, it remains effective against KV-cache eviction strategies like H2O [46]. While such lossy compression breaks the strict mathematical correspondence required for algebraic attacks, it typically preserves the semantic gist of the context. The Injection Attack successfully exploits these residual semantic vectors to hallucinate or reconstruct private data. Furthermore, the attack is highly efficient, requiring only a single generation pass. This vector underscores that a comprehensive defense must render the KV-cache semantically unintelligible, not just algebraically obfuscated.
IV Evaluation of Attacks
IV-A Experimental Setup
Models. To demonstrate attack universality, we select seven state-of-the-art LLMs spanning diverse parameter scales (1B–8B) and attention mechanisms. Our evaluation prioritizes modern GQA architectures, including the LLaMA-3.2 series (1B & 3B-Instruct), LLaMA-3.1-8B [12], and Qwen2.5-Math-7B [40]. Crucially, we incorporate DeepSeek-R1-Distill-LLaMA-8B [13] (LLaMA-3.1-8B-Distilled) to assess robustness against fine-tuned models where weight parameters diverge from their base counterparts. Finally, to verify generalizability across architectural evolutions, we evaluate legacy MHA models, specifically LLaMA-7B and LLaMA-2-7B.
Datasets. To simulate realistic privacy leakage scenarios, we utilize the LMSYS-Chat-1M dataset [48], which comprises real-world user interactions collected from inference services. We construct a test set by randomly sampling 1,000 instances, ensuring a rich variety of dialogue and instruction-following contexts. Analysis of attack generalization across domain-specific datasets is provided in Appendix -C3.
Evaluation Metrics. BERTScore [45] and ROUGE-L [23] are used to measure the similarity between the original input and the text reconstructed from the KV-cache. BERTScore, based on the all-mpnet-base-v2 model, is better at capturing semantic similarity, while ROUGE-L primarily reflects lexical-level precision and recall.
IV-B Attack Effectiveness
@@
TABLE I: Reconstruction fidelity (BERTScore and ROUGE-L) of attacks against unprotected KV-cache across different models and layers. “Collision+” denotes the attack enhanced with prior knowledge.
Type Model Metric Inversion Collision Collision+ Injection
@@
Identical LLaMA-7B BERTScore ($\uparrow$) 1.000 0.065 0.092 0.449 0.769 0.611 1.000 1.000 1.000 0.765
@@
Finetune LLaMA-3.1-8B-Distilled BERTScore ($\uparrow$) 0.083 0.062 0.081 0.642 0.492 0.635 0.894 0.258 0.762 0.610
(LLaMA-3.1-8B) ROUGE-L ($\uparrow$) 0.000 0.000 0.001 0.633 0.227 0.413 0.868 0.122 0.479 0.421
Cross-Arch LLaMA-2-7B BERTScore ($\uparrow$) 0.067 0.086 0.084 0.058 0.070 0.069 0.075 0.062 0.060 0.087
(LLaMA-7B) ROUGE-L ($\uparrow$) 0.038 0.057 0.031 0.013 0.000 0.002 0.052 0.017 0.024 0.019
This section evaluates the feasibility of reconstructing user inputs via the proposed Inversion, Collision, and Injection attacks. We systematically test these vectors against the models defined in Section IV-A. Detailed ablation studies are provided in Appendix -C.
IV-B1 Experimental Settings
Beyond standard reconstruction, we introduce two specific scenarios to evaluate attack robustness under restricted adversarial knowledge: (1) Fine-tuning Mismatch: We attack the fine-tuned LLaMA-3.1-8B-Distilled assuming the adversary only possesses weights from its base model (LLaMA-3.1-8B). This validates effectiveness when exact model weights are unavailable. (2) Cross-Architecture Mismatch: We employ LLaMA-7B parameters to attack LLaMA-2-7B to assess generalizability across disjoint architectures.
Regarding data scope, the Inversion and Collision attacks operate on single-layer KV-caches. To characterize layer-wise vulnerability, we evaluate these attacks on the first, middle, and last layers respectively. In contrast, the Injection Attack utilizes the complete, multi-layer KV-cache to leverage the model’s full semantic processing capabilities.
IV-B2 Results of KV-cache Inversion Attack
As detailed in Table I, the efficacy of the Inversion Attack generally hinges on two conditions: (1) access to the first-layer KV-cache, and (2) an algebraically invertible projection matrix, typical of MHA. Under these constraints, we achieve near-perfect reconstruction fidelity ($\approx 100\%$), whereas deeper layers yield unintelligible noise ($10\%$) due to semantic fusion. Notably, the GQA-based LLaMA-3.2-1B also exhibits high first-layer vulnerability ($\approx 100\%$). We attribute this to the model’s high projection rank relative to its hidden state dimension. This distinct property allows the least squares method to effectively recover inputs despite the non-square nature of the projection matrix.